OpenStudy (anonymous):

Solve, using linear combination. 3x + y = 4 2x + y = 5 (Points : 4) (–1, 7) (0, 4) (–3, 12) (–3, 13)

OpenStudy (anonymous):

3x+y=4 2x+y=5 [-] ------------ x = -1 put x= -1 in 3x+y=4 3*-1+y=4 -3+y=4 y=4+3 y=7 now put x= -1 and and y=7 in 2 equations to recheck the answers as they are correct or not put x = -1 and y=7 in 3x+y=4 3[-1]+7=4 -3+7=4 4=4 now put those values in 2x+y=5 2[-1]+7=5 -2+7=5 5=5 so the values of x and y are x = - 1 y = 7

OpenStudy (anonymous):

Your answer is A. @glamourglory.

OpenStudy (anonymous):

thank you guys

OpenStudy (abhisar):

Um.....i can teach you an easier way :3

OpenStudy (anonymous):

can you thank you